# Think Silently, Think Fast: Dynamic Latent Compression of LLM Reasoning Chains

### v5

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in mathematical reasoning, particularly when employing Chain-of-Thought (CoT) prompting techniques cot ; metacot . Recent advances have further highlighted the potential of this approach when combined with reinforcement learning on extended reasoning sequences o1 ; r1 ; kimi15 , revealing significant “aha-moments” in model performance. Despite these advances, a critical limitation persists: generating lengthy reasoning chains are computational costly, impeding efficiency and scalability. This inefficiency becomes particularly evident in real-world LLM applications, where extended reasoning chains create substantial server load, especially under high-concurrency conditions, underscoring the urgent need for more efficient reasoning methods.
@@
Several approaches have emerged to address this computational challenge stopsurvey ; efficientsurvey . One line of research focuses on enhancing efficiency at the token level, primarily by identifying and skipping less informative tokens tokenskip , prompting models to generate more concise reasoning steps chainfodraft ; sketchofthought , or dynamically terminating reasoning when the model exhibits high confidence in a trial answer deer . While valuable, these methods continue to operate on sparse token-based representations. A more promising direction explores reasoning within the dense latent space. Initial efforts attempt to “internalize” reasoning knowledge by curriculum learning icot-si or knowledge distillation icot-kd . Some works focus on the potential inside LLMs by looping or skipping some intermediate LLM layers ccot ; innerthinking ; loop to realize efficient reasoning. Recent innovations have introduced auto-regressive prediction of latent representations for efficient reasoning. Coconut coconut proposes to gradually replace token-level reasoning with latent representations, while CODI codi employs self-distillation to transfer CoT knowledge into latent reasoning processes. However, these methods primarily utilize fixed-length reasoning chains, resulting in suboptimal efficiency and limited adaptability. Furthermore, to the best of our knowledge, all these latent-based methods employ deterministic latent reasoning processes, overlooking the potential benefits of exploration-exploitation capability may bring about.
@@
To overcome these limitations, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses LLM reasoning chains into latent space while preserving exploration-exploitation capabilities. Our approach utilizes an auxiliary next compressed embedding prediction task in supervised fine-tuning (SFT) stage. Specifically, at each training step, CoLaR first samples a random compression factor $c\in[1,c_{\max}]$ and merges the embeddings of $c$ consecutive reasoning tokens using our Embedding Compress module. A Latent Head is then trained to predict the next compressed embeddings from the LLM’s output hidden states, which is a fully parallelized process. During inference, CoLaR is capable to auto-regressively predict dense and informative latents with the Latent Head, and automatically determine when to terminate the reasoning process with LLM’s Language Head. Rather than predicting deterministic values, the Latent Head outputs a probability distribution that produces diverse reasoning pathways for a same question input. Based on this, we further enhance CoLaR through post-training with Group Relative Policy Optimization (GRPO) reinforcement learning algorithm grpo ; dapo , which enables CoLaR to explore correct latent reasoning paths with diverse outputs and exploit those shorter ones.
@@
Our extensive evaluations on four grade-school level mathematical reasoning datasets (GSM8k gsm8k , GSM8k-hard gsmhard , SVAMP svamp , and MultiArith multiarith ) demonstrate that CoLaR achieves a $14.1\%\uparrow$ improvement in accuracy compared to state-of-the-art baseline methods at comparable compression ratios. Furthermore, CoLaR reduces reasoning chain length by $53.3\%\downarrow$ with only a $4.8\%\downarrow$ performance degradation relative to explicit CoT. Finally, experiments on a more challenging dataset MATH math demonstrates the potential of CoLaR to reinforcement learning, gaining up to $5.36\%\uparrow$ accuracy while reducing the length of reasoning chain significantly by $82.8\%\downarrow$.
@@ 2.1 Explicit LLM reasoning
Recent advances have demonstrated the strong reasoning capabilities of large language models (LLMs). The explicit reasoning approach, exemplified by Chain-of-Thought (CoT) reasoning cot , propose to prompt LLMs to generate intermediate reasoning steps through sequential token prediction before generating answers limo ; metacot ; ttr ; quietstar . Subsequent work demonstrated that reinforcement learning techniques grpo ; dapo ; vapo can further improve performance on verifiable reasoning tasks like mathematical problem-solving, revealing an “aha-moment” that significantly boosts model’s performances with longer thinking process o1 ; r1 ; kimi15 .
However, the computational cost of processing these lengthy reasoning chains remains a significant bottleneck, motivating research into efficiency optimizations. Current solutions focus on identifying and skipping redundant tokens tokenskip ; deer or encouraging more compact reasoning patterns using mathematical notations and programming language-like expressions sketchofthought ; chainfodraft . While these methods reduce reasoning chain length, they are fundamentally limited by the sequential token prediction.
@@ 2.2 Latent LLM reasoning
The first direction, knowledge internalization, aims to embed reasoning capabilities directly into the model. iCoT-SI icot-si attempts to internalize reasoning knowledge by progressively removing explicit reasoning steps during training, while Pause pause proposes training models to reason within specialized token embeddings.
The second direction exploits the hierarchical structure of transformer layers, with proposals to dynamically skip or repeat layer computations loop ; innerthinking ; ccot ; tokenassorted ; cotformer . These methods aim to reduce computational cost by selectively processing different layers.
The third direction, and most relevant to our work, explores auto-regressive latent reasoning softcot . Coconut coconut pioneered this approach by replacing token sampling with hidden state concatenation for breadth-first reasoning, while CODI codi introduces an auto-regressive latent variable model through self-distillation. However, existing methods like Coconut and CODI are limited by their reliance on fixed-length reasoning chains due to the implicit nature of latent variables. Furthermore, they employ a deterministic approach to auto-regressive latent generation, neglecting the potential for exploration-exploitation strategies to further enhance model performance, particularly within a reinforcement learning framework.
@@ 3.3 Exploration with reinforcement learning
With the Latent Head trained to predict distributions, we can sample diverse latent reasoning pathways and final answers for the same question $q$. We then apply Group Relative Policy Optimization (GRPO) algorithm grpo to reinforce correct reasoning chains and answers while penalizing incorrect ones.
@@
We simply set $r_{i}$ to $1$ when an answer is correctly predicted and to 0 otherwise. Following DAPO dapo , we remove the KL-regularization term from original GRPO implementation for efficient training.
@@ 4.1 Experimental setup
Datasets and tasks. Our method is mainly trained and evaluated on GSM8k-Aug (icot-kd, ), an augmented version of the Grade-School level Math reasoning dataset GSM8k (gsm8k, ). GSM8k-Aug comprises approximately 385k training samples and 1k test samples. We also evaluate the trained methods on three out-of-domain math reasoning datasets: (1) GSM-Hard (gsmhard, ), a modified version of GSM8K with approximately 1k test samples featuring larger magnitude numbers, (2) SVAMP (svamp, ) and (3) MultiArith multiarith , two simpler math reasoning datasets with 1k and 600 test samples, respectively. Moreover, we train and evaluate our method on a more challenging dataset MATH math , which consists of 7.5k training samples and 5k test samples, covering algebra, calculus, statistics, geometry, linear algebra and number theory. Following (coconut, ), we use two metrics: (1) Accuracy (Acc.), which measures the effectiveness of correctly predicting answers and (2) Reasoning chain length (# L), which measures efficiency by averaging the number of tokens/latents predicted in reasoning chains.
Baseline methods. We primarily compare against the following baselines: (1) CoT (cot, ), which is fine-tuned on complete reasoning chains and answers, and performs token-level reasoning before predicting answers during inference; (2) iCoT icot-si , which internalizes reasoning knowledge by gradually removing reasoning steps, and directly predicts answers during inference; (3) Coconut (coconut, ), which is fine-tuned with a curriculum process to gradually replace token-level reasoning steps with latent reasoning steps, and performs six steps of latent reasoning before predicting answers; and (4) Distill, is our reproduced version of CODI codi based on their implementation details as the code and model are not released. It self-distills token-level CoT into fixed-length latent reasoning steps, with an inference stage same to Coconut.
Implementation details. (1) Base model: unless otherwise specified, all experiments use a frozen Llama-3.2-1B-Instruct (llama3, ) backbone with a trainable LoRA module (lora, ). Following Coconut, all methods are initialized with weights from CoT-SFT to accelerate training. (2) Model checkpointing: for fair comparison, all models are trained for up to 50 epochs or 12 hours, whichever is reached first, and we choose the checkpoint that achieves the best accuracy on the validation set as the final model. (3) Hyper-parameters: we use the AdamW (adamw, ) optimizer with a fixed learning rate of 1e-4 and a weight decay of 1e-2 in SFT stage, and set the learning rate to 1e-6 in RL stage. We set $r_{max}=5$ to train CoLaR. During inference, we configure the LLM generation with a temperature of 1 and top-p of 0.9. All SFT training processes are conducted a total batch size of 256. For more implementation details, please refer to Appendix Section A.
@@ 4.4 Reinforcement learning results on the MATH dataset
We train and evaluate CoLaR with RL on the challenging MATH dataset, using two base models r1 ; qwen25 ; llama3 . The results are presented in Table 2. Our analysis of the results yields three key conclusions:
@@
(3) Base model quality impacts RL effectiveness. Supervised fine-tuning on CoT resulted in varying performance across the two base LLMs. Meanwhile, CoLaR also demonstrates a significantly larger performance gain on RL when using the higher-quality Qwen-1.5B compared to Llama-1B. This observation aligns with the findings of gandhi2025cognitive , which suggests that RL substantially activates inherent reasoning capabilities, indicating the importance of base model quality.

### v6

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in mathematical reasoning, particularly when employing Chain-of-Thought (CoT) prompting techniques Wei et al. (2022); Xiang et al. (2025). Recent advances have further highlighted the potential of this approach when combined with reinforcement learning on extended reasoning sequences Jaech et al. (2024); Guo et al. (2025); Team et al. (2025), revealing significant “aha-moments” in model performance. Despite these advances, a critical limitation persists: generating lengthy reasoning chains are computational costly, impeding efficiency and scalability. This inefficiency becomes particularly evident in real-world LLM applications, where extended reasoning chains create substantial server load, especially under high-concurrency conditions, underscoring the urgent need for more efficient reasoning methods.
@@
Several approaches have emerged to address this computational challenge Sui et al. (2025); Feng et al. (2025). One line of research focuses on enhancing efficiency at the token level, primarily by identifying and skipping less informative tokens Xia et al. (2025), prompting models to generate more concise reasoning steps Xu et al. (2025a); Aytes et al. (2025), or dynamically terminating reasoning when the model exhibits high confidence in a trial answer Yang et al. (2025). While valuable, these methods continue to operate on sparse token-based representations. A more promising direction explores reasoning within the dense latent space. Initial efforts attempt to “internalize” reasoning knowledge by curriculum learning Deng et al. (2024) or knowledge distillation Deng et al. (2023). Some works focus on the potential inside LLMs by looping or skipping some intermediate LLM layers Cheng and Van Durme (2024); Chen et al. (2025); Saunshi et al. (2025) to realize efficient reasoning. Recent innovations have introduced auto-regressive prediction of latent representations for efficient reasoning. Coconut Hao et al. (2024) proposes to gradually replace token-level reasoning with latent representations, while CODI Shen et al. (2025) employs self-distillation to transfer CoT knowledge into latent reasoning processes. However, these methods primarily utilize fixed-length reasoning chains, resulting in suboptimal efficiency and limited adaptability. Furthermore, to the best of our knowledge, all these latent-based methods employ deterministic latent reasoning processes, overlooking the potential benefits of exploration-exploitation capability may bring about.
@@
To overcome these limitations, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses LLM reasoning chains into latent space while preserving exploration-exploitation capabilities. Our approach utilizes an auxiliary next compressed embedding prediction task in supervised fine-tuning (SFT) stage. Specifically, at each training step, CoLaR first samples a random compression factor $c\in[1,c_{\max}]$ and merges the embeddings of $c$ consecutive reasoning tokens using our Embedding Compress module. A Latent Head is then trained to predict the next compressed embeddings from the LLM’s output hidden states, which is a fully parallelized process. During inference, CoLaR is capable to auto-regressively predict dense and informative latents with the Latent Head, and automatically determine when to terminate the reasoning process with LLM’s Language Head. Rather than predicting deterministic values, the Latent Head outputs a probability distribution that produces diverse reasoning pathways for a same question input. Based on this, we further enhance CoLaR through post-training with Group Relative Policy Optimization (GRPO) reinforcement learning algorithm Shao et al. (2024); Yu et al. (2025), which enables CoLaR to explore correct latent reasoning paths with diverse outputs and exploit those shorter ones.
@@
Our extensive evaluations on four grade-school level mathematical reasoning datasets (GSM8k Cobbe et al. (2021), GSM8k-hard Gao et al. (2023), SVAMP Patel et al. (2021), and MultiArith Roy and Roth (2015)) demonstrate that CoLaR achieves a $14.1\%\uparrow$ improvement in accuracy compared to state-of-the-art baseline methods at comparable compression ratios. Furthermore, CoLaR reduces reasoning chain length by $53.3\%\downarrow$ with only a $4.8\%\downarrow$ performance degradation relative to explicit CoT. Finally, experiments on a more challenging dataset MATH Hendrycks et al. (2021) demonstrates the potential of CoLaR to reinforcement learning, gaining up to $5.36\%\uparrow$ accuracy while reducing the length of reasoning chain significantly by $82.8\%\downarrow$.
@@ 2.1 Explicit LLM reasoning
Recent advances have demonstrated the strong reasoning capabilities of large language models (LLMs). The explicit reasoning approach, exemplified by Chain-of-Thought (CoT) reasoning Wei et al. (2022), propose to prompt LLMs to generate intermediate reasoning steps through sequential token prediction before generating answers Ye et al. (2025); Xiang et al. (2025); Tan et al. (2025); Zelikman et al. (2024). Subsequent work demonstrated that reinforcement learning techniques Shao et al. (2024); Yu et al. (2025); Yuan et al. (2025) can further improve performance on verifiable reasoning tasks like mathematical problem-solving, revealing an “aha-moment” that significantly boosts model’s performances with longer thinking process Jaech et al. (2024); Guo et al. (2025); Team et al. (2025).
However, the computational cost of processing these lengthy reasoning chains remains a significant bottleneck, motivating research into efficiency optimizations. Current solutions focus on identifying and skipping redundant tokens Xia et al. (2025); Yang et al. (2025) or encouraging more compact reasoning patterns using mathematical notations and programming language-like expressions Aytes et al. (2025); Xu et al. (2025a). While these methods reduce reasoning chain length, they are fundamentally limited by the sequential token prediction.
@@ 2.2 Latent LLM reasoning
The first direction, knowledge internalization, aims to embed reasoning capabilities directly into the model. iCoT-SI Deng et al. (2024) attempts to internalize reasoning knowledge by progressively removing explicit reasoning steps during training, while Pause Goyal et al. (2023) proposes training models to reason within specialized token embeddings.
The second direction exploits the hierarchical structure of transformer layers, with proposals to dynamically skip or repeat layer computations Saunshi et al. (2025); Chen et al. (2025); Cheng and Van Durme (2024); Su et al. (2025); Mohtashami et al. (2023). These methods aim to reduce computational cost by selectively processing different layers.
The third direction, and most relevant to our work, explores auto-regressive latent reasoning Xu et al. (2025b). Coconut Hao et al. (2024) pioneered this approach by replacing token sampling with hidden state concatenation for breadth-first reasoning, while CODI Shen et al. (2025) introduces an auto-regressive latent variable model through self-distillation. However, existing methods like Coconut and CODI are limited by their reliance on fixed-length reasoning chains due to the implicit nature of latent variables. Furthermore, they employ a deterministic approach to auto-regressive latent generation, neglecting the potential for exploration-exploitation strategies to further enhance model performance, particularly within a reinforcement learning framework.
@@ 3.3 Exploration with reinforcement learning
With the Latent Head trained to predict distributions, we can sample diverse latent reasoning pathways and final answers for the same question $q$. We then apply Group Relative Policy Optimization (GRPO) algorithm Shao et al. (2024) to reinforce correct reasoning chains and answers while penalizing incorrect ones.
@@
We simply set $r_{i}$ to $1$ when an answer is correctly predicted and to 0 otherwise. Following DAPO Yu et al. (2025), we remove the KL-regularization term from original GRPO implementation for efficient training.
@@ 4.1 Experimental setup
Datasets and tasks. Our method is mainly trained and evaluated on GSM8k-Aug (Deng et al., 2023), an augmented version of the Grade-School level Math reasoning dataset GSM8k (Cobbe et al., 2021). GSM8k-Aug comprises approximately 385k training samples and 1k test samples. We also evaluate the trained methods on three out-of-domain math reasoning datasets: (1) GSM-Hard (Gao et al., 2023), a modified version of GSM8K with approximately 1k test samples featuring larger magnitude numbers, (2) SVAMP (Patel et al., 2021) and (3) MultiArith Roy and Roth (2015), two simpler math reasoning datasets with 1k and 600 test samples, respectively. Moreover, we train and evaluate our method on a more challenging dataset MATH Hendrycks et al. (2021), which consists of 7.5k training samples and 5k test samples, covering algebra, calculus, statistics, geometry, linear algebra and number theory. Following (Hao et al., 2024), we use two metrics: (1) Accuracy (Acc.), which measures the effectiveness of correctly predicting answers and (2) Reasoning chain length (# L), which measures efficiency by averaging the number of tokens/latents predicted in reasoning chains.
Baseline methods. We primarily compare against the following baselines: (1) CoT (Wei et al., 2022), which is fine-tuned on complete reasoning chains and answers, and performs token-level reasoning before predicting answers during inference; (2) iCoT Deng et al. (2024), which internalizes reasoning knowledge by gradually removing reasoning steps, and directly predicts answers during inference; (3) Coconut (Hao et al., 2024), which is fine-tuned with a curriculum process to gradually replace token-level reasoning steps with latent reasoning steps, and performs six steps of latent reasoning before predicting answers; and (4) Distill, is our reproduced version of CODI Shen et al. (2025) based on their implementation details as the code and model are not released. It self-distills token-level CoT into fixed-length latent reasoning steps, with an inference stage same to Coconut.
Implementation details. (1) Base model: unless otherwise specified, all experiments use a frozen Llama-3.2-1B-Instruct (Grattafiori et al., 2024) backbone with a trainable LoRA module (Hu et al., 2022). Following Coconut, all methods are initialized with weights from CoT-SFT to accelerate training. (2) Model checkpointing: for fair comparison, all models are trained for up to 50 epochs or 12 hours, whichever is reached first, and we choose the checkpoint that achieves the best accuracy on the validation set as the final model. (3) Hyper-parameters: we use the AdamW (Loshchilov and Hutter, 2017) optimizer with a fixed learning rate of 1e-4 and a weight decay of 1e-2 in SFT stage, and set the learning rate to 1e-6 in RL stage. We set $r_{max}=5$ to train CoLaR. During inference, we configure the LLM generation with a temperature of 1 and top-p of 0.9. All SFT training processes are conducted a total batch size of 256. For more implementation details, please refer to Appendix Section A.
@@ 4.4 Reinforcement learning results on the MATH dataset
We train and evaluate CoLaR with RL on the challenging MATH dataset, using two base models Guo et al. (2025); Yang et al. (2024); Grattafiori et al. (2024). The results are presented in Table 2. Our analysis of the results yields three key conclusions:
@@
(3) Base model quality impacts RL effectiveness. Supervised fine-tuning on CoT resulted in varying performance across the two base LLMs. Meanwhile, CoLaR also demonstrates a significantly larger performance gain on RL when using the higher-quality Qwen-1.5B compared to Llama-1B. This observation aligns with the findings of Gandhi et al. (2025), which suggests that RL substantially activates inherent reasoning capabilities, indicating the importance of base model quality.